# CLARITY: Clinical Assistant for Routing, Inference, and Triage

### v1

@@ A.2.1 Insufficient Transparency in Multi-Turn Diagnostic Dialogue
One of the primary shortcomings of current medical dialogue systems is their inability to conduct structured, multi-turn reasoning. Many models generate fragmented, inconsistent, or overly simplistic responses, failing to follow a logical sequence similar to a physician’s diagnostic inquiry.
Several approaches have been proposed to address the challenges of structured diagnostic reasoning in medical dialogue systems. Chain of Diagnosis (CoD) introduces a framework designed to mimic the stepwise reasoning of physicians, systematically guiding the diagnostic process and reducing the risk of premature conclusions Chen et al. (2024). Another notable system, AMIE, leverages a self-play training mechanism within a simulated environment, enabling the model to iteratively refine both its questioning strategies and decision-making capabilities across a variety of medical scenarios Tu et al. (2024). Similarly, DISC-MedLLM enhances the depth of structured questioning by integrating medical knowledge graphs and training on real-world, multi-turn medical dialogues, ensuring that the system can navigate complex patient interactions with greater precision Bao et al. (2023).
Despite these advancements, challenges persist. Models struggle with handling ambiguous patient responses, dynamically adjusting diagnostic strategies, and effectively incorporating context across long multi-turn conversations.
@@ A.2.2 Hallucinations and Misinformation
A significant limitation of LLM-based medical dialogue systems is their tendency to generate factually incorrect or misleading information, known as hallucinations. These inaccuracies can lead to diagnostic errors and compromise patient safety.
Several strategies have been proposed to mitigate the challenges associated with medical dialogue systems. Health-LLM utilizes a retrieval-augmented generation (RAG) mechanism to ground AI-generated responses in trusted medical sources, ensuring real-time verification and reducing the risk of hallucinations Yu et al. (2025). Similarly, MEDAGENTS implements a multi-agent framework in which AI-based "experts" collaboratively cross-validate each other’s responses, thereby minimizing incorrect outputs and enhancing overall reliability Tang et al. (2024). Another approach, DISC-MedLLM, combines structured multi-turn dialogues with the integration of medical knowledge graphs to provide detailed explanations for each diagnostic step. This design not only improves the transparency of the reasoning process but also facilitates validation by enabling clinicians to trace how conclusions are reached Bao et al. (2023).
Nonetheless, hallucinations remain prevalent in complex diagnostic cases, particularly those involving rare diseases, overlapping symptoms, or mental health conditions. Ensuring real-time fact-checking and clinician-in-the-loop validation remains a pressing challenge.
@@ A.2.3 Failure to Recognize Critical Conditions
Medical AI systems often fail to detect and escalate high-risk conditions requiring urgent medical attention. This limitation reduces their effectiveness in emergency triage and high-stakes diagnostic settings.
Several approaches have been developed to improve the recognition of critical conditions in medical dialogue systems. Dual-Inf enhances diagnostic reliability by employing bidirectional inference, allowing the model to cross-reference symptoms with potential diagnoses to detect severe conditions more accurately Zhou et al. (2024b). Similarly, AMIE integrates an uncertainty-aware decision-making mechanism, which prompts the model to request additional clarifications whenever its confidence in a diagnosis is low, thereby reducing the likelihood of incorrect or premature conclusions Tu et al. (2024). Another approach, Polaris, leverages specialized risk assessment models to identify high-risk cases and escalate them for human intervention when necessary, ensuring that critical conditions receive appropriate attention Mukherjee et al. (2024).
@@ B.1 Proposed Solution to Challenges
Transparency: CLARITY ensures consistent and structured multi-turn diagnostic reasoning through FSMs, mimicking the systematic approach of experienced clinicians. In contrast, while systems like DISC-MedLLM and AMIE exhibit strong reasoning capabilities, they lack the robustness of FSM-driven workflows, leading to fragmented or inconsistent interactions in complex cases.
@@
Critical Condition Recognition: The probabilistic risk triage model in CLARITY ensures real-time identification and escalation of high-risk cases, outperforming solutions like MEDAGENTS, which lack the precision and adaptability needed for critical diagnostics.
@@ Appendix C Dialogue examples
Transparency Critical Safety Adaptability Dialogue example Doc: What’s bothering you? Pt: I have a headache. Doc: Where exactly is the pain located? Pt: The back of my head. Doc: Are you experiencing any other symptoms, such as nausea, vomiting? Pt: No. Doc: How intense is the pain? Pt: 5 out of 10. Doc: Cervicogenic headache – General practitioner. Pain in the back of the head may be related to issues in the cervical spine. Cervical osteochondrosis – Neurologist. You have a neck condition that could be causing the headache in the back of your head. Doc: Is everything clear? Feel free to ask questions! Doc: What is bothering you? Pt: High blood pressure. Doc: Do you experience chest pain or an increased heart rate? Pt: Yes. Doc: Do you have any vision problems? Pt: Yes. Doc: Your condition could be close to critical! Call 103 immediately. Wait for the response, do not hang up! Briefly explain what happened. Doc: What’s bothering you? Pt: Who are you? Doc: I am a neural network model that can assist with medical questions. How can I help you? Pt: Help me with algorithms in Python. Doc: I’m sorry, I didn’t understand your response. Could you please provide more specific information or rephrase your message? Pt: What can you help me with? Doc: I can provide consultations on health and medical issues. Please tell me more about your symptoms or concerns so that I can offer appropriate assistance. Doc: What’s bothering you? Pt: I’ve been smoking for 5 years, half a pack a day. I’ve started coughing up very thick phlegm. How can I check my lungs? Could you provide a list of examinations and doctors I need to see? Doc: You need to see a general practitioner (therapist) or a pulmonologist first. The doctor will order a chest X-ray or a CT scan of the lungs to rule out pathological changes.
@@ D.3 Readiness Estimator
The readiness estimator was trained on a corpus of 2,500 medical dialogues, with each message labeled as part of the anamnesis collection or diagnosis stage. The model used TF-IDF and contextual embeddings for feature extraction, achieving an $F_{1}$-score of 0.78. The mean absolute percentage error (MAPE) for predicting the duration of the anamnesis collection phase was 22%.
@@ Appendix I Age groups distribution
Appendix K Statistics on the recommend specialists in the expert-validated set of dialogues.
@@
where the "Text" field contains the input user message and the "OuterContex" field contains all the necessary information about the user. Inside the "OuterContex" the "Sex" and "Age" fields denote the user’s age and sex, respectively. The "UserId", "SessionId" and "ClientId" fields together form a unique dialogue identifier, which is used across the CLARITY components for dialogue context propagation.
@@
where the "Text" field contains the system output message and the "Results" field contains additional information on the dialogue context. The contents of the "Results" field depend on the dialogue state. For any dialogue state except for the Diagnosis state the "Results" field is empty. For the Diagnosis state dialog state the "Results" array contains three elements each having the following structure.

### v2

@@ A.2.1 Insufficient Transparency in Multi-Turn Diagnostic Dialogue
One of the primary shortcomings of current medical dialogue systems is their inability to conduct structured, multi-turn reasoning. Many models generate fragmented, inconsistent, or overly simplistic responses, failing to follow a logical sequence similar to a physician’s diagnostic hypotheses inquiry.
Several approaches have been proposed to address the challenges of structured diagnostic hypotheses reasoning in medical dialogue systems. Chain of Diagnosis (CoD) introduces a framework designed to mimic the stepwise reasoning of physicians, systematically guiding the diagnostic hypotheses process and reducing the risk of premature conclusions Chen et al. (2024). Another notable system, AMIE, leverages a self-play training mechanism within a simulated environment, enabling the model to iteratively refine both its questioning strategies and decision-making capabilities across a variety of medical scenarios Tu et al. (2024). Similarly, DISC-MedLLM enhances the depth of structured questioning by integrating medical knowledge graphs and training on real-world, multi-turn medical dialogues, ensuring that the system can navigate complex patient interactions with greater precision Bao et al. (2023).
Despite these advancements, challenges persist. Models struggle with handling ambiguous patient responses, dynamically adjusting diagnostic hypotheses strategies, and effectively incorporating context across long multi-turn conversations.
@@ A.2.2 Hallucinations and Misinformation
A significant limitation of LLM-based medical dialogue systems is their tendency to generate factually incorrect or misleading information, known as hallucinations. These inaccuracies can lead to diagnostic hypotheses errors and compromise patient safety.
Several strategies have been proposed to mitigate the challenges associated with medical dialogue systems. Health-LLM utilizes a retrieval-augmented generation (RAG) mechanism to ground AI-generated responses in trusted medical sources, ensuring real-time verification and reducing the risk of hallucinations Yu et al. (2025). In practice, retrieval and grounding can be supported by domain corpora that link clinical text to controlled terminologies (e.g., UMLS/ICD), such as NEREL-BIO, RuCCoN, and RuCCoD (Loukachevitch et al., 2023, 2024; Nesterov et al., 2022, 2025). Similarly, MEDAGENTS implements a multi-agent framework in which AI-based "experts" collaboratively cross-validate each other’s responses, thereby minimizing incorrect outputs and enhancing overall reliability Tang et al. (2024). Another approach, DISC-MedLLM, combines structured multi-turn dialogues with the integration of medical knowledge graphs to provide detailed explanations for each diagnostic step. This design not only improves the transparency of the reasoning process but also facilitates validation by enabling clinicians to trace how conclusions are reached Bao et al. (2023).
Nonetheless, hallucinations remain prevalent in complex diagnostic hypotheses cases, particularly those involving rare diseases, overlapping symptoms, or mental health conditions. Ensuring real-time fact-checking and clinician-in-the-loop validation remains a pressing challenge.
@@ A.2.3 Failure to Recognize Critical Conditions
Medical AI systems often fail to detect and escalate high-risk conditions requiring urgent medical attention. This limitation reduces their effectiveness in emergency triage and high-stakes diagnostic hypotheses settings.
Several approaches have been developed to improve the recognition of critical conditions in medical dialogue systems. Dual-Inf enhances diagnostic hypotheses reliability by employing bidirectional inference, allowing the model to cross-reference symptoms with potential diagnostic hypotheses to detect severe conditions more accurately Zhou et al. (2024b). Similarly, AMIE integrates an uncertainty-aware decision-making mechanism, which prompts the model to request additional clarifications whenever its confidence in a diagnostic hypotheses is low, thereby reducing the likelihood of incorrect or premature conclusions Tu et al. (2024). Another approach, Polaris, leverages specialized risk assessment models to identify high-risk cases and escalate them for human intervention when necessary, ensuring that critical conditions receive appropriate attention Mukherjee et al. (2024).
@@ B.1 Proposed Solution to Challenges
Transparency: CLARITY ensures consistent and structured multi-turn diagnostic hypotheses reasoning through FSMs, mimicking the systematic approach of experienced clinicians. In contrast, while systems like DISC-MedLLM and AMIE exhibit strong reasoning capabilities, they lack the robustness of FSM-driven workflows, leading to fragmented or inconsistent interactions in complex cases.
@@
Critical Condition Recognition: The probabilistic risk triage model in CLARITY ensures real-time identification and escalation of high-risk cases, outperforming solutions like MEDAGENTS, which lack the precision and adaptability needed for critical diagnostic hypotheses.
@@ Appendix C Dialogue examples
Transparency Critical Safety Adaptability Dialogue example Doc: What’s bothering you? Pt: I have a headache. Doc: Where exactly is the pain located? Pt: The back of my head. Doc: Are you experiencing any other symptoms, such as nausea, vomiting? Pt: No. Doc: How intense is the pain? Pt: 5 out of 10. Doc: Cervicogenic headache – General practitioner. Pain in the back of the head may be related to issues in the cervical spine. Cervical osteochondrosis – Neurologist. You have a neck condition that could be causing the headache in the back of your head. Doc: Is everything clear? Feel free to ask questions! Doc: What is bothering you? Pt: High blood pressure. Doc: Do you experience chest pain or an increased heart rate? Pt: Yes. Doc: Do you have any vision problems? Pt: Yes. Doc: Your condition could be close to critical! Call 103 immediately. Wait for the response, do not hang up! Briefly explain what happened. Doc: What’s bothering you? Pt: Who are you? Doc: I am a neural network model that can assist with medical questions. How can I help you? Pt: Help me with algorithms in Python. Doc: I’m sorry, I didn’t understand your response. Could you please provide more specific information or rephrase your message? Pt: What can you help me with? Doc: I can provide consultations on health and medical issues. Please tell me more about your symptoms or concerns so that I can offer appropriate assistance. Doc: What’s bothering you? Pt: I’ve been smoking for 5 years, half a pack a day. I’ve started coughing up very thick phlegm. How can I check my lungs? Could you provide a list of examinations and doctors I need to see? Doc: You need to see a general practitioner (therapist) or a pulmonologist first. The doctor will order a chest X-ray or a CT scan of the lungs to rule out pathological changes. Related services metrics Information collector: $P=84\,\%$, $f_{\text{rep}}=0$ Readiness estimator: $\text{MAPE}=22\,\%$ Medical-specialty selector: $R@3=96\,\%$ Emergency detector: $F_{1}=59\,\%$, FPR $2\,\%$ Moderator: $F_{1}=95\,\%$ Question detector: $F_{1}=94\,\%$
@@ D.2 Emergency Detector
Beyond the primary emergency detector, the safety stack includes:
  • A domain-specific moderator that filters unsafe topics at input and output stages.
  • Deterministic FSM gates that prevent transition to the “Diagnostic Hypothesis & Routing” state unless readiness is verified, holding low-confidence dialogues for human handoff.
  • A universal red-flag questionnaire presented to all users as a final safety net.
@@ D.3 Readiness Estimator
The readiness estimator was trained on a corpus of 2,500 medical dialogues, with each message labeled as part of the anamnesis collection or diagnostic hypotheses stage. The model used TF-IDF and contextual embeddings for feature extraction, achieving an $F_{1}$-score of 0.78. The mean absolute percentage error (MAPE) for predicting the duration of the anamnesis collection phase was 22%.
@@ D.6 Medical Specialty Selector
D.7 Language Model Details
All LLM-based components (answer generation, diagnostic hypothesis drafting, and explanation synthesis) are powered by GigaChat. We used the production configuration provided by the vendor; prompts and safety constraints are listed in Appendix D. Where relevant, we apply deterministic decoding for safety-critical paths and temperature-limited sampling for open-dialogue responses.
@@ Appendix I Age groups distribution
Appendix K Statistics on the recommended specialists in the expert-validated set of dialogues.
@@
where the "Text" field contains the input user message and the "OuterContext" field contains all the necessary information about the user. Inside the "OuterContext" the "Sex" and "Age" fields denote the user’s age and sex, respectively. The "UserId", "SessionId" and "ClientId" fields together form a unique dialogue identifier, which is used across the CLARITY components for dialogue context propagation.
@@
where the "Text" field contains the system output message and the "Results" field contains additional information on the dialogue context. The contents of the "Results" field depend on the dialogue state. For any dialogue state except for the Diagnostic hypotheses state the "Results" field is empty. For the Diagnosis state dialog state the "Results" array contains three elements each having the following structure.